MARC:

While playing mini golf, you hit the ball off of a wall to attempt at making it in the hole. The ball is located at (5,4) and the hole is located at (9,8). The point where the ball hits the wall is (7,0). Write an equation for the path of the ball and then determine if you will make your putt.

MARC:

@Zepdrix

Zepdrix:

So umm... Do they want us to come up with an absolute value function or something?

MARC:

Not sure... xD can we use the distance formula to form the eqn?

MARC:

y=\(|\frac{6}{7}x-6|\)

Zepdrix:

I'm not sure what they mean by "determine if you will make your putt.". If you do the first part that they asked for "Write an equation for the path of the ball", then you've successfully come up with an equation that touches all three points. Maybe they want a function for the ball and the wall? And then we can just let the ball miss the hole? The slope from the Ball to the Wall is \(\large\rm m=\frac{0-4}{5-7}=-2\) So we want to construct an absolute value function with slope of... oh 2 actually. Because we want a v-shape opening upwards. And we want it to pass through (7,0). \(\large\rm f(x)=2|x-7|\)

Zepdrix:

The function passes through the points (5,4) and (7,0) but ends up missing the hole by quite a bit (9,4). I hope that's what they're looking for. The instructions aren't perfectly clear.

MARC:

It's okay. Found this question from the website. xD Thanks @Zepdrix